OpenStudy (anonymous):

Solve for x in the proportion 5 over 2 times x equals 10 over the quantity 3 times x plus 3. x = 9 x = 3 x = 27 x = 2

jhonyy9 (jhonyy9):

5 10 --- = ------- cross multiplie 2x 3x +3 5(3x+3)=10(2x) divide both sides by 5 3x+3 = 2(2x) 3x+3 = 4x add to the both sides - 3x then will get x = 3

OpenStudy (anonymous):

Thankkks

jhonyy9 (jhonyy9):

was my pleasure

OpenStudy (anonymous):

Help with another, please?(:

jhonyy9 (jhonyy9):

in the question column is this exersice ?

OpenStudy (anonymous):

Haha